OpenStudy (anonymous):

In a survey of 100 outpatients who reported at a hospital one day,it was found that 70 complained of fever, 50 had stomach trouble, and 30 were injured. All the 100 patientshad one or the other of these complaints, and 44 had exactly two of them. How many patients had all three complaints?

OpenStudy (anonymous):

70+50+30=150 150-100=50 50-44=6 I think this is how you'd get the answer.

OpenStudy (anonymous):

why are we subtracting 100 from 150?

OpenStudy (anonymous):

Because there is only 100 patients

OpenStudy (anonymous):

ok